How Can i write an expression that is equivalent to 3x-(4+2x) ?
Answer by gonzo(654)

your equation:
3x-(4+2x)
remove the parentheses:
3x - 4 - 2x
combine like terms:
x - 4
your new expression that is equivalent to your original expression:
x-4 = 3x-(4+2x)
-----
to prove, let x = 0:
0-4 = 0 - (4+0)
0-4 = 0-4
-4 = -4
-----
let x = 100:
100 - 4 = 3*100 - (4 + 2*100)
96 = 300 - (204)
96 = 96
-----
equations are equivalent.